OpenStudy (anonymous):

What is the equation of the line passing through (1, 7) and parallel to the line y = –x + 5 in standard form? Answer x – y = 8 x – y = –8 x + y = 8 x + y = –8

OpenStudy (anonymous):

the third equation

OpenStudy (anonymous):

parallel means they should have the same slope \[y-7=-(x-1)\]\[y-7=-x+1\]\[y=-x+8\]\[x+y=8\]

OpenStudy (anonymous):

sarah how did you get -x? if u dont mind cud u explain it please

OpenStudy (anonymous):

Parallel lines have the same slope so both lines should have a slope of -1. We know the given line has a slope of -1 because it is written in slope-intercept form: y=mx+b===>y = –x + 5 . By the definition of slope (rise/run), and starting at the given point our line is supposed to pass through (i.e., (1,7)):\[m=\Delta y/\Delta x=\frac{y-7}{x-1}=-1\]Thus,\[y-7=-(x-1)\]etc....